# Bistability of travelling waves and wave-pinning states in a mass-conserved reaction-diffusion system:

### v3

@@ 5.2 Coexistence and bistability for moderate domain size, $L\sim\mathcal{O}(m\lambda)$
Results for domain length $L=2\lambda$ are summarized in Figs. 12 and 13. The 2WPλ solutions emerge for $s>s_{c}^{t}$ supercritically, due to finite size effects, and remain unstable until the branch terminates about $Q_{1}^{\ast}$, i.e., these solutions do not terminate at $Q_{0}^{\ast}$ like the WP solutions in large domains. The instability of WP solutions on small domains is unsurprising, cf. [64]. The 2TWλ branch is supercritical (as shown in Section 4), initially linearly stable, loses stability at $s=s_{PD}^{\lambda}$, and after a subsequent fold terminates at a parity-breaking bifurcation at $s=s_{P}^{\lambda}$. The TWs with wavelength $2\lambda$, however, are much different and resemble the structure of the EP branch. The solutions emerge from a parity-breaking bifurcation at $s=s_{P}^{2\lambda}$ of respective WP solutions, gain stability at the subsequent fold, and then lose stability at the next. Afterwards, the solutions transition to multiple peak solutions at the subsequent folds. In particular, we see that initially, the plateau of the peak expands (hollow square), then transitions into three-peak (hollow circle), and then into two-peak (hollow triangle) solutions but with different amplitudes and spacings. Finally, the branch reaches a period-doubling bifurcation (as seen with EPs) at $s=s_{PD}^{\lambda}$, where the 2TWλ branch loses stability. The WP branch resembles the bifurcation structure obtained for WP solutions in large domains: WP solutions emerge from $Q_{2}^{\ast}$ towards decreased $s$ values, and after the left fold, they gain stability, before reaching a subsequent fold and terminating at $Q_{0}^{\ast}$. Consequently, on domain size $L=2\lambda$, WP solutions stably coexist with TWs of two copies.
@@ 7 Conclusions
We analyzed a mass-conserving reaction-diffusion (RD) model Eq. 3 on a 1D domain with (mostly) periodic boundary conditions mimicking a ring of the actin cortex at the edge of a cell. Bifurcation analyses (see Sections 3, 4, and 5) reveal a rich structure of possible solutions depending not only on control parameters but also on the total mass $M$ of the $(u,v)$-subsystem. We prescribe the conditions in which steady wave-pinning (WP) solutions (a.k.a. cell polarization) can stabilize and coexist with stable travelling waves (TWs). Specifically, we study the bifurcations about a codimension-2 long wavelength (LW) and finite wavenumber Hopf (WB) instability and show how these depend on the domain length. We note, however, that the main results qualitatively persist and are not subject only to the codimension-2 LW/WB instability.
@@
The effect of mass conservation appears more profound than expected [13] (and the references therein), near local or global bifurcations. For example, the former could relate to bifurcation criticalities and spatial modulations (e.g., see Figure 19) in the unfolding of a codimension-2 LW/WB instability and also more generally in the study of a codimension-3 bifurcation, where LW, WB, and HB instabilities occur simultaneously. This construction should exist based on the results shown in Figure 3 and should extend the understanding of spatial localizations studied in the presence of HBs [29, 82]. Further studies of global bifurcations under mass-conservation could relate to the organizing centers by T-points [47] and their connection to modulated waves [44], including frequency locking for fronts exhibiting spatial oscillations (see Figure 7) and the possible existence of the Shil’nikov/WB bifurcation [36, 99]. Other avenues to further study mass-conserved RD systems and their implications for cell behaviours include the exploration of parameter space to acquire elaborated intuition about the initial conditions and perturbations that determine the selection of cell motility modes, especially informative for experimental setups, in the spirit of [38, 59, 73, 101] or the study of interactions leading to cell deformations, as inspired by [19, 20, 24, 26, 55, 61, 81, 85, 86]. Such interactions between domain geometry and reaction-diffusion dynamics promises to add new challenges and insights.
@@ Appendix B Bifurcation structure for the ventral actin wave model
Appendix C Amplitude equations derivation
We start by expanding $Q$ in powers of $\epsilon$:
where:
(30c) $F_{1}$ $=B_{LF}(\tau)e^{i(\omega_{c}t+q_{c}x)}+B_{RF}(\tau)e^{i(\omega_{c}t-q_{c}x)}+c.c.,$
and $\omega_{c},q_{c}$ denote the critical frequency and wavenumber at the instability onset. Substitution of Eq. 30 and all the scaling terms into Eq. 3 and separating into powers of $\epsilon$ (up to third order), yields at each $\mathcal{O}(\epsilon^{n})$ the corresponding equation:
@@ Appendix C Amplitude equations derivation
(34) $\textbf{\em R}_{2}=\left\{\left[\left(\gamma M-\left(3+\gamma\right)u_{c}^{\ast}\right)u_{1}+2u_{c}^{\ast}\gamma v_{1}-s_{WB}F_{1}\right]u_{1}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right),$
(35) $\textbf{\em R}_{3}=-\frac{\partial}{\partial\tau}\left(\begin{array}[]{c}u_{1}\\ v_{1}\\ F_{1}\end{array}\right)+\left\{\left[\left(2\gamma M-6u_{c}^{\ast}\right)u_{1}+2u^{\ast}_{c}\gamma\left(v_{1}-u_{1}\right)-F_{1}s_{WB}\right]u_{2}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right)\\ +\{\left[2\gamma u^{\ast}_{c}v_{2}-s_{WB}F_{2}+\left(2M\gamma-6u^{\ast}_{c}-p_{1}s_{WB}-4u^{\ast}_{c}\gamma\right)\tilde{u}^{\ast}-(p_{0}+p_{1}u^{\ast}_{c}){\widetilde{s}}\right]u_{1}\\ +2u^{\ast}_{c}\tilde{u}^{\ast}v_{1}\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right)+\left\{\left(\gamma v_{1}-u_{1}\right){u_{1}}^{2}-(s_{WB}\tilde{u}^{\ast}+{\widetilde{s}}u^{\ast}_{c})F_{1}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right).$
To avoid secular terms in $\textbf{\em R}_{n}$, we use the Fredholm alternative as a formal framework for applying the solvability condition by utilizing an inner product:
@@
$\textbf{\em N}=\left(\begin{array}[]{c}N_{1}\\ N_{2}\\ N_{3}\end{array}\right)e^{-i\left(i\omega_{c}\mp q_{c}x\right)}+c.c.$
@@
(38a) $N_{1}=$ $-\frac{\bigl{(}b+q_{c}^{2}+\gamma{u_{c}^{\ast}}^{2}+i\omega_{c}\bigr{)}\bigl{(}D_{F}q_{c}^{2}+\theta+i\omega_{c}\bigr{)}}{s_{WB}u_{c}^{\ast}\bigl{(}q_{c}^{2}+i\omega_{c}\bigr{)}}N_{3},$
(38b) $N_{2}=$ $-\frac{\bigl{(}b+\gamma{u_{c}^{\ast}}^{2}\bigr{)}\bigl{(}D_{F}q_{c}^{2}+\theta+i\omega_{c}\bigr{)}}{s_{WB}u_{c}^{\ast}\bigl{(}q_{c}^{2}+i\omega_{c}\bigr{)}}N_{3}.$
@@
$Q_{2}=\left(\begin{array}[]{c}a_{1}\\ a_{2}\\ a_{3}\end{array}\right)B_{LF}^{2}e^{2i\omega_{c}t+2iq_{c}x}+\left(\begin{array}[]{c}b_{1}\\ b_{2}\\ b_{3}\end{array}\right)B_{RF}^{2}e^{2i\omega_{c}t-2iq_{c}x}+\left(\begin{array}[]{c}c_{1}\\ c_{2}\\ c_{3}\end{array}\right)B_{LF}B_{RF}e^{2i\omega_{c}t}\\ +\left(\begin{array}[]{c}d_{1}\\ d_{2}\\ d_{3}\end{array}\right)B_{LF}\overline{B}_{RF}e^{2iq_{c}x}+E\left(\begin{array}[]{c}u_{1}\\ v_{1}\\ F_{1}\end{array}\right)+\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ h_{3}\end{array}\right)\left(|B_{LF}|^{2}+|B_{RF}|^{2}\right)+c.c.,$
where the last term is a solution of the homogeneous problem (32), $E$ is an arbitrary complex-valued constant that may be set to zero, and the other coefficients are obtained numerically using Mathematica. Finally, inserting $Q_{2}$ into the form of $\textbf{\em R}_{3}$, and expanding in powers of $e^{i\left(\pm\omega_{c}t\pm q_{c}x\right)}$, yields after applying Eq. 37 and rescaling back to original variables, the amplitude equations:
@@ D.1 Continuation and linear stability
where $Q=(u,v,F)^{\rm T}$ is the current solution, $\tilde{Q}$ is the previous solution computed along the branch, and primes denote differentiation with respect to the argument. Consequently, continuation requires three free parameters: the main bifurcation parameter, and two additional parameters for the integral constraints. For these constraints, we add two additional “auxiliary” parameters, $\kappa$ and $\psi$, which are forced to be small by the integral constraints (i.e., $\kappa,\psi<10^{-12}$ for all solutions). In AUTO, the system must be specified as a first-order ODE system. So, by defining $U:=u^{\prime}$, $V:=v^{\prime}$, and $f:=F^{\prime}$, we rewrite Eq. 3 as:

### v4

@@ 5.2 Coexistence and bistability for moderate domain size, $L\sim\mathcal{O}(m\lambda)$
Results for domain length $L=2\lambda$ are summarized in Figs. 12 and 13. The 2WPλ solutions emerge for $s>s_{c}^{t}$ supercritically, due to finite size effects, and remain unstable until the branch terminates about $Q_{1}^{\ast}$, i.e., these solutions do not terminate at $Q_{0}^{\ast}$ like the WP solutions in large domains. The instability of WP solutions on small domains is unsurprising, cf. [64]. The 2TWλ branch is supercritical (as shown in Section 4), initially linearly stable, loses stability at $s=s_{PD}^{\lambda}$, and after a subsequent fold terminates at a parity-breaking bifurcation at $s=s_{P}^{\lambda}$. The TWs with wavelength $2\lambda$, however, are much different and resemble the structure of the EP branch. The solutions emerge from a parity-breaking bifurcation at $s=s_{P}^{2\lambda}$ of respective WP solutions, gain stability at the subsequent fold, and then lose stability at the next. Afterwards, the solutions transition to multiple peak solutions at the subsequent folds. In particular, we see that initially, the plateau of the peak expands (hollow square), then transitions into three-peak (hollow circle), and then into two-peak (hollow triangle) solutions but with different amplitudes and spacings. Finally, the branch reaches a wavelength-doubling bifurcation (as seen with EPs) at $s=s_{PD}^{\lambda}$, where the 2TWλ branch loses stability. The WP branch resembles the bifurcation structure obtained for WP solutions in large domains: WP solutions emerge from $Q_{2}^{\ast}$ towards decreased $s$ values, and after the left fold, they gain stability, before reaching a subsequent fold and terminating at $Q_{0}^{\ast}$. Consequently, on domain size $L=2\lambda$, WP solutions stably coexist with TWs of two copies.
@@ 7 Conclusions
We analyzed a mass-conserving reaction-diffusion (RD) model Eq. 3 on a 1D domain with (mostly) periodic boundary conditions mimicking a ring of the actin cortex at the edge of a cell. Bifurcation analyses (see Sections 3, 4, and 5) reveal a rich structure of possible solutions depending not only on control parameters but also on the total mass $M$ of the $(u,v)$-subsystem. We prescribe the conditions in which steady wave-pinning (WP) solutions (a.k.a. cell polarization) can stabilize and coexist with stable travelling waves (TWs). Specifically, we study the bifurcations about a codimension-2 long wavelength (LW) and finite wavenumber Hopf (WB) instability and show how these depend on the domain length, i.e., unfolding via domain length, so that the main results qualitatively persist and are not subject only to the codimension-2 LW/WB instability.
@@
The effect of mass conservation appears more profound than expected [13] (and the references therein), near local or global bifurcations. For example, the former could relate to bifurcation criticalities and spatial modulations (e.g., see Figure 19) in the unfolding of a codimension-2 LW/WB instability (see Figure 5) and also more generally in the study of a codimension-3 bifurcation, where LW, WB, and HB instabilities occur simultaneously. This construction should exist based on the results shown in Figure 3 and should extend the understanding of spatial localizations studied in the presence of HBs [29, 83]. Further studies of global bifurcations under mass-conservation could relate to the organizing centers by T-points [47] and their connection to modulated waves [44], including frequency locking for fronts exhibiting spatial oscillations (see Figure 7) and the possible existence of the Shil’nikov/WB bifurcation [36, 100]. Other avenues to further study mass-conserved RD systems and their implications for cell behaviours include the exploration of parameter space to acquire elaborated intuition about the initial conditions and perturbations that determine the selection of cell motility modes, especially informative for experimental setups, in the spirit of [38, 59, 74, 102] or the study of interactions leading to cell deformations, as inspired by [19, 20, 24, 26, 55, 61, 82, 86, 87]. Such interactions between domain geometry and reaction-diffusion dynamics promises to add new challenges and insights.
@@ Appendix B Bifurcation structure for the ventral actin wave model
Appendix C Derivation of amplitude equations
Here we discuss the main algebraic steps in the derivation of the amplitude equations (17). Substitution of Eq. 16, with $s=s_{WB}+\epsilon^{2}\widetilde{s}$ and $Q^{\ast}=Q_{c}^{\ast}(s_{WB})+\epsilon^{2}\widetilde{Q}^{\ast}$, into Eq. 3, and separating into powers of $\epsilon$ (up to third order), yields at each $\mathcal{O}(\epsilon^{n})$ the corresponding equation:
@@ Appendix C Derivation of amplitude equations
(34) $\textbf{R}_{2}=\left\{\left[\left(\gamma M-\left(3+\gamma\right)u_{c}^{\ast}\right)u_{1}+2u_{c}^{\ast}\gamma v_{1}-s_{WB}F_{1}\right]u_{1}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right),$
(35) $\textbf{R}_{3}=-\frac{\partial}{\partial\tau}\left(\begin{array}[]{c}u_{1}\\ v_{1}\\ F_{1}\end{array}\right)+\left\{\left[\left(2\gamma M-6u_{c}^{\ast}\right)u_{1}+2u^{\ast}_{c}\gamma\left(v_{1}-u_{1}\right)-F_{1}s_{WB}\right]u_{2}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right)\\ +\{\left[2\gamma u^{\ast}_{c}v_{2}-s_{WB}F_{2}+\left(2M\gamma-6u^{\ast}_{c}-p_{1}s_{WB}-4u^{\ast}_{c}\gamma\right)\tilde{u}^{\ast}-(p_{0}+p_{1}u^{\ast}_{c}){\widetilde{s}}\right]u_{1}\\ +2u^{\ast}_{c}\tilde{u}^{\ast}v_{1}\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right)+\left\{\left(\gamma v_{1}-u_{1}\right){u_{1}}^{2}-(s_{WB}\tilde{u}^{\ast}+{\widetilde{s}}u^{\ast}_{c})F_{1}\right\}\left(\begin{array}[]{c}1\\ -1\\ 0\end{array}\right).$
To avoid secular terms in $\textbf{R}_{n}$, we use the Fredholm alternative as a formal framework for applying the solvability condition by utilizing an inner product:
@@
[…]
where $\varrho=1+p_{0}s_{WB}+p_{1}s_{WB}u_{c}^{*}+3u_{c}^{*2}+2\gamma u_{c}^{*2}-2Mu_{c}^{*}\gamma$, and the components of the nullvector
$\textbf{N}=\left(\begin{array}[]{c}N_{1}\\ N_{2}\\ N_{3}\end{array}\right)e^{i\left(\omega_{c}t\pm q_{c}x\right)}+c.c.$
@@
(38a) $N_{1}=$ $-\frac{\bigl(b+q_{c}^{2}+\gamma{u_{c}^{\ast}}^{2}-i\omega_{c}\bigr)\bigl(D_{F}q_{c}^{2}+\theta-i\omega_{c}\bigr)}{s_{WB}u_{c}^{\ast}\bigl(q_{c}^{2}-i\omega_{c}\bigr)}N_{3},$
(38b) $N_{2}=$ $-\frac{\bigl(b+\gamma{u_{c}^{\ast}}^{2}\bigr)\bigl(D_{F}q_{c}^{2}+\theta-i\omega_{c}\bigr)}{s_{WB}u_{c}^{\ast}\bigl(q_{c}^{2}-i\omega_{c}\bigr)}N_{3}.$
@@
$Q_{2}=\left(\begin{array}[]{c}a_{1}\\ a_{2}\\ a_{3}\end{array}\right)B_{LF}^{2}e^{2i\omega_{c}t+2iq_{c}x}+\left(\begin{array}[]{c}b_{1}\\ b_{2}\\ b_{3}\end{array}\right)B_{RF}^{2}e^{2i\omega_{c}t-2iq_{c}x}+\left(\begin{array}[]{c}c_{1}\\ c_{2}\\ c_{3}\end{array}\right)B_{LF}B_{RF}e^{2i\omega_{c}t}\\ +\left(\begin{array}[]{c}d_{1}\\ d_{2}\\ d_{3}\end{array}\right)B_{LF}\overline{B}_{RF}e^{2iq_{c}x}+\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ h_{3}\end{array}\right)\left(|B_{LF}|^{2}+|B_{RF}|^{2}\right)+E\left(\begin{array}[]{c}u_{1}\\ v_{1}\\ F_{1}\end{array}\right)+c.c.,$
where the last term is a solution of the homogeneous problem (32), $E$ is an arbitrary complex-valued constant that is set to zero, and the other coefficients are obtained numerically using Mathematica by solving 31 for respective terms at order $n=2$. Finally, inserting $Q_{2}$ into the form of $\textbf{R}_{3}$, and expanding in powers of $e^{i\left(\pm\omega_{c}t\pm q_{c}x\right)}$, yields after applying Eq. 37 and rescaling back to original variables, the amplitude equations:
@@ D.1 Continuation and linear stability
where $Q=(u,v,F)^{\rm T}$ is the current solution, $\tilde{Q}$ is the previous solution computed along the branch, and primes denote differentiation with respect to the argument. Consequently, continuation requires three free parameters: the main bifurcation parameter and two additional parameters for the integral constraints. For these constraints, we add two additional “auxiliary” parameters, $\kappa$ and $\psi$, which are forced to be small by the integral constraints (i.e., $\kappa,\psi<10^{-12}$ for all solutions). In AUTO, the system must be specified as a first-order ODE system. So, by defining $U:=u^{\prime}$, $V:=v^{\prime}$, and $f:=F^{\prime}$, we rewrite Eq. 3 as: